How many molecules of water are in .895 mol of H2O
1
Expert's answer
2020-03-30T10:58:49-0400

Solution:

n(H2O) = 0.895 mol

1 mole has 6.022×1023 molecules.

Thus, 0.895 moles has:

0.895 * 6.022×1023 = 5.38969×1023 = 5.39×1023 molecules of water.


Answer: 5.39×1023 molecules of water are in 0.895 mol of H2O.